3logx=log512-log8
Solve and check.

Answer by lwsshak3(11628)

3logx=log512-log8
Solve and check.
3logx=log512-log8
3logx-log512+log8=0
log[(x^3)*8)/512]=0
convert to exponential form
10^0=(x^3)*8)/512=1
8x^3=512
Take cube root of both sides
2x=8
x=4
Check:
3log(4)≈1.8062
log512-log8≈1.0862
left side=right side